OpenStudy (theghostlycreeper):

What is the value of n? −6.5(n−5)=18.2

OpenStudy (studygurl14):

First distribute the -6.5

KjSaif (kjsaif):

Step 1: Simplify both sides of the equation. −6.5(n−5)=18.2 Simplify: (Show steps) −6.5n+32.5=18.2 Step 2: Subtract 32.5 from both sides. −6.5n+32.5−32.5=18.2−32.5 −6.5n=−14.3 Step 3: Divide both sides by -6.5. −6.5n/−6.5=−14.3/−6.5 n=2.2

KjSaif (kjsaif):

Answer is n = 2.2 :)